Can you find the x-intercepts for the parabola defined by the equation below? And how do you do it im so confused?

y = 3x^2 + 18x + 15

Answer by scott8148(6628)

the x-intercepts are when y equals zero (also called roots)

0 = 3x^2 + 18x + 15

dividing by 3 ___ 0 = x^2 + 6x + 5

factoring ___ 0 = (x + 5)(x + 1)

x + 5 = 0 ___ x = -5

x + 1 = 0 ___ x = -1